# Heterogeneous RBCs via deep multi-agent reinforcement learning

### v1

@@ Abstract.
Current macroeconomic models with agent heterogeneity can be broadly divided into two main groups. Heterogeneous-agent general equilibrium (GE) models, such as those based on Heterogeneous Agents New Keynesian (HANK) or Krusell-Smith (KS) approaches, rely on GE and ‘rational expectations’, somewhat unrealistic assumptions that make the models very computationally cumbersome, which in turn limits the amount of heterogeneity that can be modelled. In contrast, agent-based models (ABMs) can flexibly encompass a large number of arbitrarily heterogeneous agents, but typically require the specification of explicit behavioural rules, which can lead to a lengthy trial-and-error model-development process. To address these limitations, we introduce MARL-BC, a framework that integrates deep multi-agent reinforcement learning (MARL) with Real Business Cycle (RBC) models. We demonstrate that MARL-BC can: (1) recover textbook RBC results when using a single agent; (2) recover the results of the mean-field KS model using a large number of identical agents; and (3) effectively simulate rich heterogeneity among agents, a hard task for traditional GE approaches. Our framework can be thought of as an ABM if used with a variety of heterogeneous interacting agents, and can reproduce GE results in limit cases. As such, it is a step towards a synthesis of these often opposed modelling paradigms.
@@
multi-agent systems, reinforcement learning, real business cycle
@@ 1. Introduction
Macroeconomic modelling has traditionally relied on General Equilibrium (GE) models employing representative agents, such as the Real Business Cycle (RBC) and New Keynesian (NK) models. However, a well-known limitation of representative-agent models is their inability to account for agent heterogeneity Kirman (1992); Stoker (1993).
To capture heterogeneity, two main modelling approaches emerged in macroeconomics: (i) heterogeneous-agent GE models, and (ii) Agent-Based Models (ABMs). The first approach embeds heterogeneity directly into a GE framework, typically extending a representative agent model through a ‘mean-field’ approach. A mean-field approach, in this context, implies that individual decisions are computed in the presence of the effect of aggregate quantities, which play a role similar to a mean field in the physics literature. Influential early examples aimed at introducing heterogeneous households into an otherwise standard RBC setting are those of Aiyagari (1994) and of Krusell and Smith (1998). In recent years, Heterogeneous Agent New Keynesian (HANK) models have incorporated similar extensions into NK models Kaplan et al. (2018); Auclert et al. (2025). These GE heterogeneous-agent models come at a high cost, as each agent is assumed to have ‘rational expectations’, meaning that, in practice, it needs to keep track of the entire wealth or income distribution as state variables to make any decision, an assumption that can be thought of as both unrealistic and too computationally burdensome Moll (2025). Furthermore, the high computational costs involved significantly restrict the degree of heterogeneity achievable with GE models. For example, often heterogeneity appears ‘ex-post’, meaning that all agents are initially identical and differentiate only as a result of the individual random shocks they undergo during the simulation. The second major approach is agent-based modelling. ABMs abandon the representative-agent and rational expectation assumptions entirely. They simulate economies ‘from the bottom up’ with many interacting agents that are both ex-ante and ex-post heterogeneous, have no rational expectations, and instead are endowed with boundedly rational heuristic decision rules Farmer and Foley (2009); Dawid and Gatti (2018); Farmer (2024); Axtell and Farmer (2025); Borsos et al. (2025). This flexibility of ABMs is also the feature for which they face the strongest criticism. Because in ABMs the modeller typically needs to decide directly the agents’ behavioural rules, it can be difficult to correctly navigate the arbitrariness in their specification and pin down realistic rules Windrum et al. (2007); Benedetti et al. (2022).
Only very recently, researchers have begun exploring Reinforcement Learning (RL) Sutton et al. (1998) – and especially Multi-Agent Reinforcement Learning (MARL) Albrecht et al. (2024) – as a novel way to model heterogeneous agents in macroeconomics Atashbar and Aruhan Shi (2022); Cook and Palmer (2025). RL agents learn optimal (or nearly optimal) behaviours through repeated interaction with an environment to maximise their reward signals over time. The RL learning paradigm seems to offer a natural synthesis between the extremes of GEs and ABMs: agents can be boundedly rational and diverse, yet their behaviours emerge endogenously from a principled optimisation process (learning to maximise rewards), so that the modeller can avoid searching directly for a behavioural rule Brusatin et al. (2024). Moreover, modern deep RL techniques can solve complex, high-dimensional and nonlinear optimisation problems, which would be difficult to tackle with traditional methods.
@@
Figure 1. The MARL-BC framework and two limit cases. The left orange square contains a schematic illustration of the MARL-BC framework. A population of $n$ heterogeneous RL household agents, $i=1,\dots,n$, possesses distinct productivities of capital $\kappa_{i}$ and of labour $\lambda_{i}$, and provides capitals $k^{i}_{t}$ and labours $\ell^{i}_{t}$. These individual inputs aggregate into total capital $K_{t}$ and total labour $L_{t}$, which drive production $Y_{t}$ via a Cobb-Douglas function. In the markets, the individual interest rates $r^{i}_{t}$ and the wages $w^{i}_{t}$ are assumed to be proportional to the corresponding individual productivities. Wealth $a^{i}_{t}$ is given by past depreciated capital $(1-\delta)k^{i}_{t}$, earned wages $w^{i}_{t}\ell^{i}_{t}$, and earned interest $r^{i}_{t}k^{i}_{t}$. Each agent needs to decide the fraction $\hat{c}^{i}_{t}$ of wealth to consume and the amount of labour to provide. Consumption and labour go into the instantaneous reward $R^{i}_{t}$ of the agent, which the agent learns to maximise in the long run with a discount factor $\beta$. The two right orange boxes represent classic macroeconomic models as limiting cases of our framework. The RBC model (top) is obtained using a representative agent ($n=1$) and an AR(1) process for the technology shocks $A_{t}$. The Krusell–Smith mean-field model (bottom) is obtained using many ex-ante identical agents and technology and labour switching between discrete values.
@@
  • We develop the Multi Agent Real Business Cycle (MARL-BC) framework, a MARL-based framework that extends the classic RBC model with multiple households characterised by rich and flexible heterogeneity.
  • We demonstrate that training of MARL-BC is computationally feasible, employing state-of-the-art RL algorithms such as Proximal Policy Optimisation (PPO) or Soft Actor Critic (SAC) or Deep Deterministic Policy Gradient (DDPG).
@@
  • We further show that MARL-BC can recover the mean-field Krusell-Smith model when using a large population of ex-ante identical agents.
@@
  • Finally, we illustrate the capability of the framework in simulating a richer heterogeneity among agents, impossible to achieve with a GE approach.
@@
Our results demonstrate the value and relevance of the MARL-BC framework for a growing community of researchers in both economics and computer science, thus paving the way for numerous applications and future investigations. Accordingly, we made the code for the MARL-BC easy-to-use and available in open source at https://github.com/fedegabriele/MARL-BC.
@@
Related work. The idea of using RL as a model for realistic human decision making is attracting increasing research and development interest for economic and financial simulations Tilbury (2022); Atashbar and Aruhan Shi (2022). Influential early contributions that sparked interest in this direction were the studies in Johanson et al. (2022) and Zheng et al. (2022), where deep multi-agent RL was employed to simulate emergent economic behaviours within simplified toy economies. Initially, its adoption occurred mostly within the financial sector, where RL was successfully applied to model various trading strategies Ardon et al. (2021); Mascioli et al. (2024); Gao et al. (2023); Evans and Ganesh (2024), and has led to the development of specialised open-source platforms Amrouni et al. (2021); Ardon et al. (2023). In macroeconomics, recent literature has begun exploring RL techniques to extend classic GE frameworks Cook and Palmer (2025). For instance, Curry et al. (2023) and Mi et al. (2024) employ MARL for tax policy optimisation within MARL macroeconomic environments and Dwarakanath et al. (2024) employ MARL to optimise tax credit strategies. In Chen et al. (2025) and Atashbar and Aruhan Shi (2023), RL is used to solve representative-agent GE models. In Hill et al. (2021), RL is leveraged to solve a hybrid model that combines features of both macro ABMs and general equilibrium models. Parallel studies focused on extending RL with ABMs. For instance, in Brusatin et al. (2024), the authors incorporate RL firms within a standard macroeconomic ABM, while in Glielmo et al. (2023) the authors use RL to enhance the calibration of macro ABMs. In Agrawal et al. (2025), RL is used to design robust macroeconomic policies resilient to model misspecification and uncertainty. Similarly, Evans et al. (2025) proposes a general framework combining adaptive agent behaviours through RL within ABMs. Our work is also related to another line of ongoing research which aims to tackle the curse of dimensionality of GE models with deep learning, but without using RL (Fernández-Villaverde et al., 2024; Fernández-Villaverde, 2025). Finally, our work can contribute to research studying effects of bounded rationality, typically done without RL, in existing ABMs or GE models Dosi et al. (2020); Asano et al. (2021).
@@
Significance. Our work stands out clearly in comparison with the described literature for the following reasons. Existing literature at the economics end of the spectrum has mostly focussed on single-agent RL, showing that it can recover the policy functions of representative‑agent GE models. However, research in economics has not explored how to build more expressive models using a multi-agent RL approach. Conversely, the computer‑science literature has experimented with multi-agent RL, showing that the approach can give rise to rich emergent economic behaviours. However, research in computer science has mostly overlooked even the most fundamental models from macroeconomics and it has not convincingly shown how multi-agent RL can extend such models. In our view, this generated a clear knowledge gap between the two disciplines, which currently impedes communication and mutually beneficial exchanges of ideas. For example, this implied that, on the one hand, economic modelling could not leverage the full potential of multi-agent RL as a viable strategy for extending its traditional modelling frameworks and, on the other hand, that research and development in computer science could not progress rapidly towards the construction of frameworks of actual relevance for the economics discipline.
@@
Our work represents a clear step forward in bridging this knowledge gap, as it provides a foundation for linking research in two disciplines. For this reason, we believe it has a significant potential for stimulating fruitful mutual exchanges. Specifically, we bridge the two research lines by proposing a framework that (i) exactly reproduces the representative‑agent RBC solution when $n=1$, (ii) converges to the KS mean‑field equilibrium as $n\gg 1$, and (iii) extends beyond both limits to easily accommodate a high degree of discrete heterogeneity that GE methods cannot easily handle.
@@
The rest of this work is structured as follows. In Sec. 2 we describe the MARL-BC framework. In Sec. 3, we provide details on the training and testing procedure we followed for the numerical experiments. In Sec. 4, we describe the results of our numerical experiments, dividing the discussion into three parts: the replication of textbook RBC results, the replication of mean-field Krusell-Smith results, and the illustration of the full capabilities of our framework. Finally, in Sec. 5, we conclude.
@@ 2. The MARL-BC framework
We here describe the Multi-Agent Reinforcement Learning Business Cycle (MARL-BC) framework, in its general form. This will allow us to analyse the traditional models, such as the textbook RBCs or the Krusell-Smith, simply as limiting cases. Our framework and the mentioned limiting cases are illustrated in Figure 1. Specifically, the orange box in the figure contains the general MARL-BC framework. The framework is composed of two pieces: the heterogeneous RBC environment, depicted as the four solid blue boxes, is described in Sec. 2.1, while the RL household problem, depicted as the dashed blue box, is described in Sec. 2.2.
@@ 2.1. The heterogeneous RBC environment
The model consists of $n$ types of households $i=1,\dots,n$ and a single firm. At each time $t$, each type of household (which we will also refer to as a ‘household’) can own a different amount of capital $k^{i}_{t}$ and can provide a different amount of labour $\ell^{i}_{t}$. Furthermore, each agent has a different productivity of capital $\kappa^{i}$ and of labour $\lambda^{i}$. We here assume that capital and labour productivities are fixed in time, but the model can be easily extended to allow them to vary.
@@
$K_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\kappa^{i}\,k^{i}_{t},$ $\qquad L_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\lambda^{i}\,$ (1)
which are averages of the households’ variables, weighted by the corresponding productivities. Although these are formally written averages, we take the population mass to be one (a convention that also ensures consistency with standard GE models), so averages coincide with aggregates, quantities are expressed in a per capita basis, and the firm’s conditions in Eq. (3) apply.
@@
with elasticity to output $\alpha$, and technology $A_{t}$.
@@ 2.2. The RL households
Action space. The RL households’ action at each time step $t$ is a tuple $(\hat{c}^{i}_{t},\ell^{i}_{t})$ of two continuous numbers, the consumption fraction $\hat{c}^{i}_{t}$ and the labour $\ell^{i}_{t}$. While theoretically both of these fall in the range $(0,1)$, for numerical stability we clip them in the range $(0.01,0.99)$.
The consumption fraction $\hat{c}^{i}_{t}$ determines $c^{i}_{t}$, which is the amount of household’s wealth spent on consumption
@@
By choosing $\hat{c}^{i}_{t}$ and $\ell^{i}_{t}$, the households need to balance current consumption and leisure with future earnings, as will be made clearer in the rest of this section.
@@
$\mathbf{x}^{i}_{t}=(k^{i}_{t},K_{t},\ell^{i}_{t-1},L_{t-1},A_{t},\kappa^{i},\lambda^{i}),$ (7)
containing respectively the values of individual and aggregate capital, past individual and aggregate labour, and technology. In practice, not all of these variables are needed to obtain interesting dynamics or to reproduce traditional macroeconomic models, and, in fact, we will sometimes select a subset as specified in the results section.
@@
$\pi^{i}:\mathbf{x}^{i}_{t}\rightarrow(\hat{c}^{i}_{t},\ell^{i}_{t})$ (8)
giving consumption fraction $\hat{c}^{i}_{t}$ and labour $\ell^{i}_{t}$ as a function of the state of the household and of the economy $\mathbf{x}^{i}_{t}$. The policy $\pi^{i}$ is learned by maximising $\mathcal{R}^{i}$, the expected sum of discounted rewards
@@
It increases with increasing consumption $c^{i}_{t}$ and leisure $(1-\ell^{i}_{t})$, with $b>0$ controlling the trade-off between the two. Rewards are known as ‘utilities’ in the economic literature. Furthermore, the specific function in Eq. (10) is very commonly used since the logarithmic utility is, in fact, a special case of the well-known ‘Constant Relative Risk Aversion’ (CRRA) utility function when the coefficient of relative risk aversion equals one Romer (2018).
Game theoretic considerations. From a game-theoretic perspective, the households can be thought to participate in a stochastic game with $n$ players $i=1,\dots,n$, a common state, individual actions $(\hat{c}_{t}^{i},\ell_{t}^{i})$, a transition kernel induced by Eqs. (14), and per-period payoffs $R^{i}_{t}$. This, in turn, ensures the existence of optimal policies for the households assuming a solution concept such as Markov perfect equilibrium Maskin and Tirole (2001). However, the corresponding problem is well known to be computationally hard Daskalakis et al. (2009). Consequently, although the approach would be theoretically possible with a full specification of states, actions, transitions and rewards, it is impractical due to the high complexity and dimensionality of the underlying model and its strategy space. In addition, relying on the full information set of the model is an unrealistic assumption, as real-world agents typically lack full observability of their environment. In contrast, we train independent learners each having access only to the partial information set $\mathbf{x}_{t}^{i}$ and optimising for a policy $\pi^{i}$ that approximates a best response to the evolving behaviour of the others.
@@
$\delta$ Capital depreciation $\{1,0.025\}\mid 0.025\mid 0.25$
@@
Table 1. Parameters used in the experiments. List of parameters used in the numerical experiments. When three values are reported, separated by a vertical bar, they refer respectively to three result categories (RBC, KS and General) as specified in the main text. For the AR(1) technology process we use $\rho=0.9$ and $\sigma=0.01$. The KS technology process is described in the main text.

### v2

@@ Abstract.
Current macroeconomic models with agent heterogeneity can be broadly divided into two main groups. Heterogeneous-agent general equilibrium (GE) models, such as those based on Heterogeneous Agent New Keynesian (HANK) or Krusell–Smith (KS) approaches, rely on GE and ‘rational expectations’, somewhat unrealistic assumptions that make the models very computationally cumbersome, which in turn limits the amount of heterogeneity that can be modelled. In contrast, agent-based models (ABMs) can flexibly encompass a large number of arbitrarily heterogeneous agents, but typically require the specification of explicit behavioural rules, which can lead to a lengthy trial-and-error model-development process. To address these limitations, we introduce MARL-BC, a framework that integrates deep multi-agent reinforcement learning (MARL) with real business cycle (RBC) models. We demonstrate that MARL-BC can: (1) recover textbook RBC results when using a single agent; (2) recover the results of the mean-field KS model using a large number of identical agents; and (3) effectively simulate rich heterogeneity among agents, a hard task for traditional GE approaches. Our framework can be thought of as an ABM if used with a variety of heterogeneous interacting agents, and can reproduce GE results in limit cases. As such, it is a step towards a synthesis of these often opposed modelling paradigms.
@@
multi-agent systems; reinforcement learning; real business cycle
@@ 1. Introduction
Figure 1. The MARL-BC framework and two limit cases. The left orange square contains a schematic illustration of the MARL-BC framework. A population of $n$ heterogeneous RL households, $i=1,\dots,n$, possesses distinct productivities of capital $\kappa_{i}$ and of labour $\lambda_{i}$, and provides capital $k^{i}_{t}$ and labour $\ell^{i}_{t}$. These individual inputs aggregate into total capital $K_{t}$ and total labour $L_{t}$, which drive production $Y_{t}$ via a Cobb-Douglas function. The individual interest rates $r^{i}_{t}$ and the wages $w^{i}_{t}$ are assumed to be proportional to the corresponding individual productivities. Wealth $a^{i}_{t}$ is given by past depreciated capital $(1-\delta)k^{i}_{t}$, earned wages $w^{i}_{t}\ell^{i}_{t}$, and earned interest $r^{i}_{t}k^{i}_{t}$. Each agent decides the fraction $\hat{c}^{i}_{t}$ of wealth to consume and the amount of labour to provide. These choices determine the instantaneous reward $R^{i}_{t}$, which the agent learns to maximise in the long run. The two right orange boxes represent classic macroeconomic models as limiting cases of our framework. The RBC model (top) is obtained using a representative agent ($n=1$) and an AR(1) process for the technology shocks $A_{t}$. The Krusell–Smith mean-field model (bottom) is obtained using many ex-ante identical agents and technology and labour switching between discrete values.
Macroeconomic modelling has traditionally relied on General Equilibrium (GE) models employing representative agents, such as the real business cycle (RBC) and new Keynesian (NK) models. However, a well-known limitation of representative-agent models is their inability to account for agent heterogeneity Kirman1992; stoker1993empirical.
To capture heterogeneity, two main modelling approaches emerged in macroeconomics: (i) heterogeneous-agent GE models, and (ii) agent-based models (ABMs). The first approach embeds heterogeneity directly into a GE framework, typically extending a representative agent model through a ‘mean-field’ approach. In this setting, individuals act given aggregate quantities, which act as a ‘mean field’, in analogy with physics. Influential early examples aimed at introducing heterogeneous households into the RBC setting are those of Aiyagari1994 and of KrusellSmith1998. More recently, Heterogeneous Agent New Keynesian (HANK) models have incorporated similar extensions into NK models KaplanMollViolante2018; auclert2025fiscal. These GE heterogeneous-agent models come at a high cost, as each agent is assumed to have ‘rational expectations’ and needs to keep track of the entire wealth or income distribution as state variable to make any decision, an assumption that can be thought of as both unrealistic and too computationally burdensome moll2024trouble. Furthermore, the high computational costs involved significantly restrict the degree of heterogeneity achievable with GE models. For example, often heterogeneity appears ‘ex-post’, meaning that all agents are initially identical and differentiate only as a result of the individual shocks they undergo during the simulation. The second major approach is agent-based modelling. ABMs abandon the representative-agent and rational expectations assumptions entirely. They simulate economies ‘from the bottom up’ with many interacting agents that are both ex-ante and ex-post heterogeneous, have no rational expectations, and instead are endowed with boundedly rational heuristic decision rules FarmerFoley2009; dawid2018agent; farmer2024making; axtell2025agent; borsos2025agent. This flexibility of ABMs is also the reason why they face the strongest criticism. Because in ABMs the modeller typically needs to decide the agents’ behaviour directly, it can be difficult to correctly navigate the arbitrariness in their specification and pin down realistic rules windrum2007empirical; benedetti2022black.
Only very recently, researchers have begun exploring reinforcement learning (RL) sutton1998reinforcement – and especially multi-agent reinforcement learning (MARL) albrecht2024multi – as a novel way to model heterogeneous agents in macroeconomics atashbar2022deep; cook2025reinforcement. RL agents learn optimal (or nearly optimal) behaviours through repeated interaction with an environment to maximise their reward signals over time. The RL learning paradigm seems to offer a natural synthesis between the extremes of GE models and ABMs: agents can be boundedly rational and diverse, yet their behaviours emerge endogenously from a principled optimisation process (learning to maximise rewards), so that the modeller can avoid searching directly for a behavioural rule brusatin2024simulating. Moreover, modern deep RL techniques can solve complex, high-dimensional and nonlinear optimisation problems, which would be difficult to tackle with traditional methods.
@@
  • We develop the Multi-Agent Reinforcement Learning Business Cycle (MARL-BC), a MARL-based framework that extends the classic RBC model with multiple households characterised by rich and flexible heterogeneity.
  • We demonstrate that the training of MARL-BC is computationally feasible, employing state-of-the-art RL algorithms such as Proximal Policy Optimisation (PPO) or Soft Actor Critic (SAC) or Deep Deterministic Policy Gradient (DDPG).
@@
  • We further show that MARL-BC can recover the mean-field Krusell–Smith model when using a large population of ex-ante identical agents.
@@
  • Finally, we illustrate the capability of the framework to simulate richer heterogeneity among agents, impossible to achieve with a GE approach.
@@
Our results demonstrate the value and relevance of the MARL-BC framework for a growing community of researchers in both economics and computer science, thus paving the way for numerous applications and future investigations.
@@
Related work. The idea of using RL as a model for realistic human decision making is attracting increasing research and development interest for economic and financial simulations tilbury2022reinforcement; atashbar2022deep. Influential early contributions that sparked interest in this direction were the studies in johanson2022emergent and zheng2022ai, where deep multi-agent RL was employed to simulate emergent economic behaviours within simplified toy economies. Initially, its adoption occurred mostly within the financial sector, where RL was successfully applied to model trading strategies in different markets ardon2021towards; mascioli2024financial; gao2023deeper; Evans2024; bacaloni2025natural, and has led to the development of specialised open-source platforms amrouni2021abides; phantom2023. In macroeconomics, recent literature has begun exploring RL techniques to extend classic GE frameworks cook2025reinforcement. For instance, chen2021deep; hinterlang2021optimal use RL for monetary policy, while CurryEtAl2023; curry2022analyzing; MiEtAl2023; dwarakanath2024tax employ MARL for tax policy optimisation. In AtashbarShi2023 and yang2025structural, RL is used to solve representative-agent and heterogeneous GE models respectively, while in hill2021solving it is leveraged to solve a hybrid model that combines features of both macro ABMs and GE models. Parallel studies focused on extending RL with ABMs. For instance, in brusatin2024simulating, the authors incorporate RL firms within a standard ABM, while in glielmo2023reinforcement the authors use RL to enhance the calibration of ABMs. In agrawal2025robust, RL is used to design robust macroeconomic policies resilient to model misspecification and uncertainty. Similarly, Evans2025 proposes a general framework combining adaptive agent behaviours through RL within ABMs. Our work is also related to another line of ongoing research which aims to tackle the curse of dimensionality of GE models with deep learning, but without using RL (fernandez2024taming; fernandez2025deep). Finally, our work can contribute to research studying the effects of bounded rationality, typically done without RL, in existing ABMs or GE models dosi2020rational; asano2021emergent.
@@
Significance. Our work stands out clearly in comparison with the described literature for the following reasons. Existing literature at the economics end of the spectrum has mostly focused on single-agent RL, showing that it can solve representative‑agent GE models. However, research in economics has not explored how to build more expressive models using a multi-agent RL approach. Conversely, the computer‑science literature has experimented with multi-agent RL, showing that it can give rise to rich emergent economic behaviours. However, research in computer science has mostly overlooked fundamental models from macroeconomics, and it has not convincingly shown how multi-agent RL can extend such models. In our view, this has created a clear knowledge gap between the two disciplines, which currently impedes communication and mutually beneficial exchanges of ideas. For example, this implied that, on the one hand, economic modelling could not leverage the full potential of multi-agent RL as a viable strategy for extending its traditional modelling frameworks and, on the other hand, that research and development in computer science could not progress rapidly towards the construction of frameworks of actual relevance for the economics discipline.
@@
Our work represents a clear step forward in bridging this knowledge gap, as it provides a foundation for linking research in two disciplines. For this reason, we believe it has significant potential for stimulating fruitful mutual exchanges. Specifically, we bridge the two research lines by proposing a framework that (i) exactly reproduces the representative‑agent RBC solution when $n=1$, (ii) converges to the KS mean‑field equilibrium as $n\gg 1$, and (iii) extends beyond both limits to easily accommodate a high degree of discrete heterogeneity that GE methods cannot easily handle.
@@
The rest of this work is structured as follows. In Sec. 2 we describe the MARL-BC framework. In Sec. 3, we provide details on the training and testing procedure we followed for the numerical experiments. In Sec. 4, we describe the results of our numerical experiments, dividing the discussion into three parts: the replication of textbook RBC results, the replication of mean-field Krusell–Smith results, and the illustration of the full capabilities of our framework. Finally, in Sec. 5, we conclude.
@@ 2. The MARL-BC framework
We describe here the Multi-Agent Reinforcement Learning Business Cycle (MARL-BC) framework, in its general form. This will allow us to analyse the traditional models, such as the textbook RBCs or the Krusell–Smith, simply as limiting cases. Our framework and the mentioned limiting cases are illustrated in Figure 1. Specifically, the orange box in the figure contains the general MARL-BC framework. The framework is composed of two pieces: the heterogeneous RBC environment, depicted as the four solid blue boxes, is described in Sec. 2.1, while the RL household problem, depicted as the dashed blue box, is described in Sec. 2.2.
@@ 2.1. The heterogeneous RBC environment
The model consists of $n$ types of households $i=1,\dots,n$ and a single firm. At each time $t$, each type of household (which we will also refer to as a ‘household’) can own a different amount of capital $k^{i}_{t}$ and can provide a different amount of labour $\ell^{i}_{t}$. Furthermore, each agent has a different productivity of capital $\kappa^{i}$ and of labour $\lambda^{i}$. We assume here that capital and labour productivities are fixed in time, but the model can be easily extended to allow them to vary.
@@
$K_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\kappa^{i}\,k^{i}_{t},$ $\qquad L_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\lambda^{i}\,\ell^{i}_{t},$ (1)
which are averages of the households’ variables, weighted by the corresponding productivities. Although these are formally written averages, we take the population mass to be one (a convention that also ensures consistency with standard GE models), so averages coincide with aggregates, quantities are expressed on a per capita basis, and the firm’s conditions in Eq. (3) apply.
@@
with output elasticity of capital $\alpha$, and technology $A_{t}$.
@@ 2.2. The RL households
Action space. The RL households’ action at each time step $t$ is a tuple $(\hat{c}^{i}_{t},\ell^{i}_{t+1})$ of two continuous numbers, the consumption fraction $\hat{c}^{i}_{t}$ and the labour $\ell^{i}_{t+1}$. While theoretically both of these fall in the range $(0,1)$, for numerical stability we clip them in the range $(0.01,0.99)$.
The consumption fraction $\hat{c}^{i}_{t}$ determines $c^{i}_{t}$, which is the amount of the household’s wealth spent on consumption
@@
By choosing $\hat{c}^{i}_{t}$ and $\ell^{i}_{t+1}$, the households need to balance current consumption and leisure with future earnings, as will be made clearer in the rest of this section.
@@
$\mathbf{x}^{i}_{t}=(k^{i}_{t},K_{t},\ell^{i}_{t},L_{t},A_{t},\kappa^{i},\lambda^{i}),$ (7)
containing the values of individual and aggregate capital, past individual and aggregate labour, and technology. In practice, not all of these variables are needed to obtain interesting dynamics or to reproduce traditional macroeconomic models, and, in fact, we will sometimes select a subset as specified in the results section.
@@
$\pi^{i}:\mathbf{x}^{i}_{t}\rightarrow(\hat{c}^{i}_{t},\ell^{i}_{t+1})$ (8)
giving consumption fraction $\hat{c}^{i}_{t}$ and labour $\ell^{i}_{t+1}$ as a function of the state of the household and of the economy $\mathbf{x}^{i}_{t}$. The policy $\pi^{i}$ is learned by maximising $\mathcal{R}^{i}$, the expected sum of discounted rewards
@@
It increases with increasing consumption $c^{i}_{t}$ and leisure $(1-\ell^{i}_{t})$, with $b>0$ controlling the trade-off between the two. Rewards are known as ‘utilities’ in the economic literature. Furthermore, the specific function in Eq. (10) is very commonly used since the logarithmic utility is, in fact, a special case of the well-known ‘Constant Relative Risk Aversion’ (CRRA) utility function when the coefficient of relative risk aversion equals one romer2018advanced.
Game-theoretic considerations. From a game-theoretic perspective, the households can be thought of as participating in a stochastic game with $n$ players $i=1,\dots,n$, a common state, individual actions $(\hat{c}_{t}^{i},\ell_{t+1}^{i})$, a transition kernel induced by Eqs. (14), and per-period payoffs $R^{i}_{t}$. This, in turn, ensures the existence of optimal policies for the households assuming a solution concept such as Markov perfect equilibrium maskin2001markov. However, the corresponding problem is well known to be computationally hard daskalakis2009complexity. Consequently, although the approach would be theoretically possible, it is impractical due to the high complexity and dimensionality of the underlying model and its strategy space. In addition, relying on the full information set of the model is an unrealistic assumption, as real-world agents typically lack full observability of their environment. In contrast, we train independent learners each having access only to the partial information set $\mathbf{x}_{t}^{i}$ and optimising for a policy $\pi^{i}$ that approximates a best response to the evolving behaviour of the others. In our numerical experiments, we show that for sufficiently large $n$, the resulting MARL scheme reproduces the equilibrium behaviour of KrusellSmith1998. However, from a theoretical standpoint, establishing convergence of independent MARL dynamics to mean field equilibria in stochastic games remains a challenging and open problem yongacoglu2024mean; yardim2023policy.
@@
$\delta$ Capital depreciation $\{1,0.025\}\mid 0.025\mid 0.025$
@@
Table 1. Parameters used in the experiments. When three values are reported, separated by a vertical bar, they refer respectively to the three result categories (RBC, KS, and General) as specified in the main text. For the AR(1) technology process we use $\rho=0.9$ and $\sigma=0.01$. The KS technology process is described in the main text.